A man in a speed boat which could only travel at 50km/h was driving his boat up a river which flowed with a constant speed. As he was driving his spare life jacket fell of the boat, 20 minutes later the man noticed the jacket was gone and turned back to get it. He found it 3km away from where he lost it. How fast was the river flowing?
Answer by venugopalramana(3286)

A man in a speed boat which could only travel at 50km/h was driving his boat up a river which flowed with a constant speed(S SAY).
RELATIVE SPEED WHILE TRAVELLING UPSTREAM = 50-S KMPH
As he was driving his spare life jacket fell of the boat (AT P SAY), 20 minutes
DISTANCE TRAVELLED BY MAN IN 20 MTS=1/3 HR.UP STREAM = (50-S)/3 KM =PQ SAY
DISTANCE TRAVELLED BY JACKET IN 20 MTS DOWN STREAM = S/3 KM =PR SAY
HENCE INTIAL DISTANCE OF SEPERATION= D =PQ+PR= S/3 + (50-S)/3 = 50/3 KM.
later the man noticed the jacket was gone and turned back to get it.
TIME TAKEN TO GET IT = D / RELATIVE SPEED TRAVELLING DOWN STREAM BETWEEN MAN AND JACKET = D/(50+S-S)=D/50
=50/(3*50) =1/3 HRS.
DISTANCE TRAVELLED BY HIM IN THIS TIME = QT = (50+S)/3
He found it 3km away from where he lost it.
PT=3 KM = QT-PQ = [(50+S)/3]-[(50-S)/3]
50+S-50+S=3*3=9
2S=9
S=4.5 KMPH
How fast was the river flowing?